OpenStudy (anonymous):

Find the equation of the circle tangent to the x-axis at (6,0) and tangent to the y-axis

OpenStudy (anonymous):

@amistre64 @phi

OpenStudy (amistre64):

your thoughts?

OpenStudy (anonymous):

the tangent line goes through (6,0) but I'm not sure about the y axis part

OpenStudy (amistre64):

well, we can always solve for k i spose ... unless you can determine a useful property of a circle and its tangents

OpenStudy (anonymous):

how do we solve for K... ?

OpenStudy (amistre64):

(x-h)^2 + (y-k)^2 = r^2 is the equation of a circle centered at h,k right?

OpenStudy (anonymous):

so (6-h)^2+(0-k)^2=r^2 ?

OpenStudy (amistre64):

that is one way to try an approach ...

OpenStudy (amistre64):



OpenStudy (amistre64):

tangent lines have normals that point to the center .

OpenStudy (amistre64):

to the center of a circle tha tis

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok how do we start this question ..im a bit confuse

OpenStudy (amistre64):

depend on what you know .... my approach is not necessarily going to be your approach.

OpenStudy (amistre64):

if we are tangent to the x axis, at 6,0 ... what does that limit our radius to if we also have to be tangent to the y axis?

OpenStudy (anonymous):

r=3?

OpenStudy (amistre64):

not what im thinking ...

OpenStudy (amistre64):

will our center be aong the x=6 axis?

OpenStudy (anonymous):

idk can it?

OpenStudy (amistre64):

well, if normals to the tangents point to the center ... and we are tangent to the x axis at x=6,